OpenStudy (anonymous):

Someone please help me this stuff is so frustrating:/ Question 20 (Essay Worth 5 points) [10.07] Brian is creating a collage on a piece of cardboard that has an area of 160r4 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of the square root of r to the seventh power square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

OpenStudy (kirbykirby):

\(160r^4\)?

OpenStudy (anonymous):

It already gives you that. It says make an equation tho ughhhh I am so confused.

OpenStudy (kirbykirby):

I am just making sure this is what is meant. 160r4 doesn't show me that it's an exponent because of the formatting

OpenStudy (anonymous):

Yeah its an exponent

OpenStudy (kirbykirby):

each piece has area \((\sqrt{r})^7=r^{\frac{1}{2}*7}=r^{7/2}\)

OpenStudy (kirbykirby):

to know the number of pieces that fit, just divide the total area but the are of each piece: # of pieces = \[\frac{160r^4}{r^{7/2}}\]

OpenStudy (kirbykirby):

\[=160r^{1/2}\]

OpenStudy (anonymous):

where'd you get 1/2?

OpenStudy (kirbykirby):

\[4-\frac{1}{2}=\frac{8}{2}-\frac{7}{2}=\frac{1}{2}\] This is related to the law of exponents that says \[\frac{a^M}{a^N}=a^{M-N}\]

OpenStudy (kirbykirby):

oops sorry I meant 4-\frac{7}{2}=\frac{8}{2}-\frac{7}{2}=\frac{1}{2}

OpenStudy (anonymous):

I am still confused. You have to break it down more because I am such a blonde lol:)

OpenStudy (kirbykirby):

\[4-\frac{7}{2}=\frac{8}{2}-\frac{7}{2}=\frac{1}{2}\]

OpenStudy (kirbykirby):